# Access Controls Will Solve the Dual-Use Dilemma

### v3

@@ 1 Introduction
In this paper, we argue that informative, hard-to-fabricate real-world context can be obtained through user-level verifications such as ID checks, institutional affiliations, or government-issued certifications. We address the dual-use dilemma with two contributions. As our primary contribution, we show how this verified context can be used jointly with content analysis in a safety framework based on access controls (Lampson, 1974). In the framework, model outputs are classified into content categories, and the system verifies whether the user has the required credentials to access the detected category. We also describe how the framework addresses both issues caused by the dual-use dilemma: over-refusals and under-refusals. Additionally, we propose a novel theoretical approach to content category classification based on recent methods for robust unlearning, UNDO (Lee et al., 2025) and gradient routing (Cloud et al., 2024). This approach avoids the capability gap between a model and its monitors that can make output monitoring methods non-robust (Jin et al., 2024).
Current approaches to AI safety force crude trade-offs between blanket restrictions that stifle beneficial uses and permissive policies that enable misuse. Our framework represents a first step toward solving the challenge of “detection and authorization of dual-use capability at inference time” highlighted by a recent survey of problems in technical AI governance (Reuel et al., 2025). By giving model providers better tools for contextual safety decisions, access control frameworks could benefit all stakeholders — enabling users to access more capabilities whilst providing regulators with a way to avoid blunt regulatory instruments.
@@ 2 Current Safety Methods Don’t Solve the Dual-Use Dilemma
The dual-use dilemma causes two issues: over-refusals and under-refusals. Over-refusals reduce model utility for legitimate users, which is clearly undesirable. Under-refusals are equally problematic because they enable decomposition attacks (Glukhov et al., 2023, 2024). These attacks transform clearly harmful queries, such as “How to modify a virus to avoid immune detection?”, into series of mundane grey-zone questions, such as the question about viral surface proteins from Figure 1. Safety systems would refuse the harmful query but do not refuse the grey-zone ones. Through these attacks, adversaries exploit under-refusals in a way that compromises system safety.
@@ 2.1 Unlearning: Non-Contextual Removal of Concepts
Unlearning methods aim to remove specific knowledge, concepts, or capabilities from a model after training (Liu et al., 2024). Their goal is to eliminate the model’s ability to generate harmful content while preserving other capabilities.
Unlearning faces significant technical challenges even for preventing behaviours that are clearly harmful. As noted by Cooper et al. (2024) and Barez et al. (2025), capabilities are hard to define, hard to remove without side effects, and hard to trace back to specific data points. Furthermore, many unlearning approaches mask rather than truly remove the targeted knowledge (Deeb & Roger, 2025).
@@ 2.2 Safety Training: The Model Reacts to Context
Safety training methods modify the model’s training process to align its outputs with human preferences. This category includes safety pre-training (Maini et al., 2025), RLHF (Christiano et al., 2023), and safety finetuning.
Unlike unlearning, these methods are contextual. They don’t remove capabilities entirely but train the model to selectively deploy them based on, among other things, the perceived legitimacy and harmlessness of the request. However, these qualities are entirely inferred from user-supplied information, such as the request text or chat history. Without access to trustworthy real-world context, the model cannot make truly informed decisions about grey-zone scenarios. For example, models are susceptible to attacks that fabricate in-chat context (Zeng et al., 2024), or attacks that diminish models’ sensitivity to it (Russinovich et al., 2025). While modifying these methods to incorporate external contextual information is possible in theory, it would likely result in a more opaque and less modular system than making similar modifications to post-processing methods.
@@ 2.3 Post-Processing: External Systems React to Context
Post-processing methods are systems that classify user inputs and system outputs for the purposes of steering the underlying model, or monitoring and filtering its responses. Sometimes, these methods are used for usage monitoring, as is the case with Anthropic’s Clio (Tamkin et al., 2024; Handa et al., 2025), other times, they are used for safety, as with Llama Guard (Inan et al., 2023) and Constitutional Classifiers (Sharma et al., 2025). However, similarly to safety training, the “real-world” context these methods work with is currently inferred mostly from user-supplied content and thus untrustworthy and vulnerable to attacks, as evidenced by the many jailbreaks that successfully target current production systems (Zhang et al., 2025). Nevertheless, these methods could be modified to incorporate external contextual information, potentially serving as a foundation for more trustworthy, contextual safety mechanisms. We discuss this option in Section 3.4.
@@ 3.2 Content Categories
Model providers will develop content categories by adapting existing risk frameworks with the help of domain experts. In biology, experts could build on biosafety levels (BSL) (Centers for Disease Control and Prevention & National Institutes of Health, 2020) and dual-use research of concern policies (United States Government, 2012). However, since existing frameworks typically categorize only high-level concepts like organisms or compounds, experts would need to decompose them into smaller, more specific components suitable for knowledge access control.
@@
Evidence from chemistry suggests this approach is at least sometimes feasible: the risk schedules of the Chemical Weapons Convention already identify not just controlled compounds but also their precursors (Organisation for the Prohibition of Chemical Weapons, 1993), demonstrating successful decomposition into components. Nevertheless, some harmful applications might not decompose so neatly; we discuss this limitation in Section 4.
@@ 3.3 Verification Mechanisms
For moderate-risk categories, model providers could use established identity verification services like Stripe Identity (Stripe, Inc., 2024) or institutional systems like ORCID (ORCID, Inc., 2024). These systems provide global, standardized, low-friction solutions with one-time costs under \$2 per user. They would serve primarily to maintain audit trails for post-incident investigation and provide a deterrent effect.
@@ 3.4 Implementing Content Classification
To implement this, we could concentrate distributed knowledge into specialized expert modules using a novel method that would combine UNDO (Lee et al., 2025) and gradient routing (Cloud et al., 2024). The approach would first unlearn knowledge belonging to content categories from the base model, then distil this unlearned model into a new model with expert modules for each category. During distillation, gradients from examples in each content category would be routed exclusively through their associated expert modules, while the model would be trained to activate experts only when generating relevant content.
@@
Moving from theoretical approaches to more practical ones, post-processing methods such as Constitutional Classifiers (Sharma et al., 2025) offer a proven approach to content classification. They operate independently of the model, allowing for rapid deployment and iteration, and they could be adapted to detect content categories and trigger checks of user verifications. However, for latency reasons, there is sometimes a capability gap between the model and the post-processing system, which adversaries can exploit to evade detection (Jin et al., 2024; Kumar et al., 2025).
@@ 4.2 Domain-specific Approaches to Under-Refusals
Models contain a lot of safety-relevant knowledge about pathogen biology; a recent study found that frontier models have more tacit knowledge than 94% of expert virologists (Götting et al., 2025). Consider an access control system so badly calibrated that all pathogen-related queries would require verification. This represents an upper bound on friction since real implementations would be far more targeted. First, we note that this system would deter at least some adversaries from performing decomposition attacks, improving on the safety of the status quo. At the same time, the impact on utility is minimal: only 0.85% of queries would experience added friction (estimated based on the Anthropic Economic Index, see Appendix A for details). This would likely be acceptable to model providers; for context, existing safety systems like Constitutional Classifiers incorrectly refuse around 0.5% queries (Sharma et al., 2025).
@@ 4.4 Limitations and Open Questions
Since the system links real-world identities to requests, it poses risks of surveillance and information leakage. Model providers should establish clear privacy policies and use privacy-preserving systems like Clio (Tamkin et al., 2024) for logging and auditing.
@@ Acknowledgements
We thank Jakub Kryś, and Dennis Akar for their feedback on a draft of this paper. We thank Joseph Miller, Alex Cloud, Alex Turner, and Jacob Goldman-Wetzler for discussions on gradient routing.
@@ Appendix A Estimating the Number of Requests Related to the Biology of Pathogens
To estimate how many user requests are related to the biology of pathogens, we used the second version of the Anthropic Economic Index (Handa et al., 2025), a dataset of 1 million anonymized conversations from the Free and Pro tiers of Claude.ai. In the dataset, the conversations are clustered by topic, and the proportion of each topic in the whole dataset is given. For example, the topic “Help with agricultural business, research, and technology projects” makes up 0.15% of the requests in the dataset. There are three levels of topic granularity; we use the lowest, most granular level.

### v4

@@ 1 Introduction
In this paper, we argue that informative, hard-to-fabricate real-world context can be obtained through user-level verifications such as ID checks, institutional affiliations, or government-issued certifications. We address the dual-use dilemma with two contributions. As our primary contribution, we show how this verified context can be used jointly with content analysis in a safety framework based on access controls (butler1974). In the framework, model outputs are classified into content categories, and the system verifies whether the user has the required credentials to access the detected category. We also describe how the framework addresses both issues caused by the dual-use dilemma: over-refusals and under-refusals. Additionally, we propose a novel theoretical approach to content category classification based on recent methods for robust unlearning, UNDO (lee2025distillationrobustifiesunlearning) and gradient routing (cloud2024gradientroutingmaskinggradients). This approach avoids the capability gap between a model and its monitors that can make output monitoring methods non-robust (jin2024jailbreakinglargelanguagemodels).
Current approaches to AI safety force crude trade-offs between blanket restrictions that stifle beneficial uses and permissive policies that enable misuse. Our framework represents a first step toward solving the challenge of “detection and authorization of dual-use capability at inference time” highlighted by a recent survey of problems in technical AI governance (reuel2025openproblemstechnicalai). By giving model providers better tools for contextual safety decisions, access control frameworks could benefit all stakeholders — enabling users to access more capabilities whilst providing regulators with a way to avoid blunt regulatory instruments.
@@ 2 Current Safety Methods Don’t Solve the Dual-Use Dilemma
The dual-use dilemma causes two issues: over-refusals and under-refusals. Over-refusals reduce model utility for legitimate users, which is clearly undesirable. Under-refusals are equally problematic because they enable decomposition attacks (glukhov2023llmcensorshipmachinelearning; glukhov2024breachthousandleaksunsafe). These attacks transform clearly harmful queries, such as “How to modify a virus to avoid immune detection?”, into series of mundane grey-zone questions, such as the question about viral surface proteins from Figure 1. Safety systems would refuse the harmful query but do not refuse the grey-zone ones. Through these attacks, adversaries exploit under-refusals in a way that compromises system safety.
@@ 2.1 Unlearning: Non-Contextual Removal of Concepts
Unlearning methods aim to remove specific knowledge, concepts, or capabilities from a model after training (liu2024rethinkingmachineunlearninglarge). Their goal is to eliminate the model’s ability to generate harmful content while preserving other capabilities.
Unlearning faces significant technical challenges even for preventing behaviours that are clearly harmful. As noted by cooper2024machineunlearningdoesntthink and barez2025openproblemsmachineunlearning, capabilities are hard to define, hard to remove without side effects, and hard to trace back to specific data points. Furthermore, many unlearning approaches mask rather than truly remove the targeted knowledge (deeb2025unlearningmethodsremoveinformation).
@@ 2.2 Safety Training: The Model Reacts to Context
Safety training methods modify the model’s training process to align its outputs with human preferences. This category includes safety pre-training (maini2025safetypretraininggenerationsafe), RLHF (christiano2023deepreinforcementlearninghuman), and safety finetuning.
Unlike unlearning, these methods are contextual. They don’t remove capabilities entirely but train the model to selectively deploy them based on, among other things, the perceived legitimacy and harmlessness of the request. However, these qualities are entirely inferred from user-supplied information, such as the request text or chat history. Without access to trustworthy real-world context, the model cannot make truly informed decisions about grey-zone scenarios. For example, models are susceptible to attacks that fabricate in-chat context (zeng2024johnnypersuadellmsjailbreak), or attacks that diminish models’ sensitivity to it (russinovich2025greatwritearticlethat). While modifying these methods to incorporate external contextual information is possible in theory, it would likely result in a more opaque and less modular system than making similar modifications to post-processing methods.
@@ 2.3 Post-Processing: External Systems React to Context
Post-processing methods are systems that classify user inputs and system outputs for the purposes of steering the underlying model, or monitoring and filtering its responses. Sometimes, these methods are used for usage monitoring, as is the case with Anthropic’s Clio (tamkin2024clioprivacypreservinginsightsrealworld; handa2025economictasksperformedai), other times, they are used for safety, as with Llama Guard (inan2023llamaguardllmbasedinputoutput) and Constitutional Classifiers (sharma2025constitutionalclassifiersdefendinguniversal). However, similarly to safety training, the “real-world” context these methods work with is currently inferred mostly from user-supplied content and thus untrustworthy and vulnerable to attacks, as evidenced by the many jailbreaks that successfully target current production systems (zhang2025outputconstraintsattacksurface). Nevertheless, these methods could be modified to incorporate external contextual information, potentially serving as a foundation for more trustworthy, contextual safety mechanisms. We discuss this option in Section 3.4.
@@ 3.2 Content Categories
Model providers will develop content categories by adapting existing risk frameworks with the help of domain experts. In biology, experts could build on biosafety levels (BSL) (CDC_BMBL_2020) and dual-use research of concern policies (USG_DURC_2012). However, since existing frameworks typically categorize only high-level concepts like organisms or compounds, experts would need to decompose them into smaller, more specific components suitable for knowledge access control.
@@
Evidence from chemistry suggests this approach is at least sometimes feasible: the risk schedules of the Chemical Weapons Convention already identify not just controlled compounds but also their precursors (OPCW_CWC_1993), demonstrating successful decomposition into components. Nevertheless, some harmful applications might not decompose so neatly; we discuss this limitation in Section 4.
@@ 3.3 Verification Mechanisms
For moderate-risk categories, model providers could use established identity verification services like Stripe Identity (stripe_identity_2024) or institutional systems like ORCID (orcid_2024). These systems provide global, standardized, low-friction solutions with one-time costs under \$2 per user. They would serve primarily to maintain audit trails for post-incident investigation and provide a deterrent effect.
@@ 3.4 Implementing Content Classification
To implement this, we could concentrate distributed knowledge into specialized expert modules using a novel method that would combine UNDO (lee2025distillationrobustifiesunlearning) and gradient routing (cloud2024gradientroutingmaskinggradients). The approach would first unlearn knowledge belonging to content categories from the base model, then distil this unlearned model into a new model with expert modules for each category. During distillation, gradients from examples in each content category would be routed exclusively through their associated expert modules, while the model would be trained to activate experts only when generating relevant content. Concurrent work adopted and successfully evaluated a similar approach for controlling access to information the model learned through fine-tuning (jayaraman2025permissionedllmsenforcingaccess). Our proposed approach is more general and could be applied to knowledge obtained in any training stage.
@@
Moving from theoretical approaches to more practical ones, post-processing methods such as Constitutional Classifiers (sharma2025constitutionalclassifiersdefendinguniversal) offer a proven approach to content classification. They operate independently of the model, allowing for rapid deployment and iteration, and they could be adapted to detect content categories and trigger checks of user verifications. However, for latency reasons, there is sometimes a capability gap between the model and the post-processing system, which adversaries can exploit to evade detection (jin2024jailbreakinglargelanguagemodels; kumar2025freelunchguardrails).
@@ 4.2 Domain-specific Approaches to Under-Refusals
Models contain a lot of safety-relevant knowledge about pathogen biology; a recent study found that frontier models have more tacit knowledge than 94% of expert virologists (gotting2025virologycapabilitiestestvct). Consider an access control system so badly calibrated that all pathogen-related queries would require verification. This represents an upper bound on friction since real implementations would be far more targeted. First, we note that this system would deter at least some adversaries from performing decomposition attacks, improving on the safety of the status quo. At the same time, the impact on utility is minimal: only 0.85% of queries would experience added friction (estimated based on the Anthropic Economic Index, see Appendix A for details). This would likely be acceptable to model providers; for context, existing safety systems like Constitutional Classifiers incorrectly refuse around 0.5% queries (sharma2025constitutionalclassifiersdefendinguniversal).
@@ 4.4 Limitations and Open Questions
Since the system links real-world identities to requests, it poses risks of surveillance and information leakage. Model providers should establish clear privacy policies and use privacy-preserving systems like Clio (tamkin2024clioprivacypreservinginsightsrealworld) for logging and auditing.
@@ Acknowledgements
We thank Jakub Kryś, and Dennis Akar for their feedback on a draft of this paper. We thank Joseph Miller, Alex Cloud, Alex Turner, and Jacob Goldman-Wetzler for discussions on gradient routing. We are grateful to Ryan Kidd and the ICML Technical AI Governance Workshop for funding support that enabled the presentation of this work.
@@ Appendix A Estimating the Number of Requests Related to the Biology of Pathogens
To estimate how many user requests are related to the biology of pathogens, we used the second version of the Anthropic Economic Index (handa2025economictasksperformedai), a dataset of 1 million anonymized conversations from the Free and Pro tiers of Claude.ai. In the dataset, the conversations are clustered by topic, and the proportion of each topic in the whole dataset is given. For example, the topic “Help with agricultural business, research, and technology projects” makes up 0.15% of the requests in the dataset. There are three levels of topic granularity; we use the lowest, most granular level.